# A Comprehensive Survey on Knowledge Distillation

### v1

@@ IV-B3 Adversarial Robust Knowledge Distillation
As previously discussed, KD transfers knowledge from a large, pre-trained teacher network to a smaller student network, yet it does not inherently guarantee robustness against adversarial attacks. Adversarial Robust Distillation (ARD) [166] addresses this limitation by combining adversarial training with distillation. In ARD, the objective is to minimize the discrepancy between the student’s predictions and the teacher’s predictions under adversarial perturbations, defined as follows
@@
Building upon ARD, several approaches have been developed to enhance the transfer of robust knowledge. One group refines teacher prediction reliability: RSLAD [167] replaces the classification loss in eq. 6 with a KL loss between teacher and student predictions to capitalize on robust soft labels from adversarially trained teachers, while introspective adversarial distillation (IAD)[168] and MTARD[169] further allow selective trust in teacher outputs and employ dual teachers for clean and adversarial scenarios, respectively. Another group enhances alignment and adaptability by synchronizing gradients and adaptively deriving perturbations through IGDM [170] and AdaAD [171]; additional refinements are achieved in SmaraAD [172] and IGKD-BML [173] via Spearman Correlation, Class Activation Mapping, and attention-guided distillation to better emulate the teacher’s decision-making process.
@@
Advanced strategies tackle limitations in robustness transfer by introducing novel training schemes. DARWIN [174] incorporates intermediate adversarial samples along with a triplet-based loss to balance natural, adversarial, and intermediate distributions, while PeerAiD [175] trains a peer model alongside the student for dynamic adaptation. Furthermore, DGAD [176] partitions the dataset into standard and adversarial distillation groups with consistency regularization to manage data imbalance, and both STARSHIP [177] and TALD [178] further bolster robustness by transferring statistical attributes and sampling diverse adversarial examples via a Teacher Adversarial Local Distribution using Stein Variational Gradient Descent. These grouped techniques collectively advance the robustness and overall performance of student models in adversarial environments.
IV-C Multi-teacher Distillation
@@
In contrast to typical distillation scenarios where the student is trained using a single teacher, multi-teacher algorithms aim to train the student network by amalgamating knowledge from multiple teachers. This approach has the potential to enhance the student’s performance by leveraging an ensemble of teachers with diverse knowledge, thereby bolstering the student’s generalization capabilities. This concept was initially proposed in [10], which utilized the average of logits from multiple teachers as softened labels. Subsequent works have explored variations on this theme: [44] employed a voting strategy, while [28] recommended perturbing the teacher’s logits by adding noise, treating these perturbed outputs as distinct teachers that can act as a regularizer. [179] utilized a pool of teachers, randomly selecting a teacher for distillation each time. [180] trained multiple teachers, with each focusing on a subset of the dataset, and employed a voting system to distill the knowledge of each expert teacher. [99] adopted a step-by-step strategy in which the student, at each stage, served as a teacher for the student in the subsequent step.
@@
In recent years, a variety of multi-teacher methods have been introduced, each addressing distinct issues. Managing the knowledge aggregation adaptively poses a challenge due to the capacity gap between the student and each teacher [181, 182, 183, 184, 185, 186]. [184] employs adaptive temperature and a diverse aggregation strategy to enhance distillation performance, while [185] separates the vanilla KD loss and assigns adaptive weights to each teacher based on the entropy of their predictions. [186] leverages data relation knowledge to dynamically allocate weights to teachers, and [182] employs a meta-network for weighting each teacher effectively.
@@
Other methodologies have been proposed to effectively aggregate the knowledge of teachers for distillation [187, 188, 189, 190, 191]. [188] trains two distinct teachers, one for distilling intricate features and another for transferring general decision features. [189] and [190] introduce additional branches to the student network for learning the features of teachers, while [191] suggests a progressive approach for distilling knowledge from multiple teachers. Another significant research avenue in multi-teacher distillation involves using ensemble pre-trained teachers, referred to as Knowledge Amalgamation (KA) [192, 193, 194, 195, 196, 197, 198, 199]. Through KA, publicly available trained networks can be repurposed in the context of multi-teacher distillation. Some approaches involve employing multiple teachers, each trained on different datasets [174] or tasks [192, 187]. [193] initially extracts task-specific knowledge from heterogeneous teachers sharing the same sub-task and then combines this extracted knowledge to construct the student network. [199] proposes a KA framework based on uncertainty suppression, and [194] suggests training a unified classifier from pre-trained classifiers from each source along with an unlabeled set of generic data.
@@
IV-D Cross-modal Distillation
In many scenarios, some modalities have rich annotations, while others lack sufficient labels. Cross-Modality Knowledge Distillation (CMKD) addresses this challenge by leveraging a well-annotated modality to improve the learning of a less-annotated one. This technique enables the model to benefit from the additional modality during training while remaining functional, even when that modality is unavailable at inference time. In 2016, [200] introduced an innovative framework to transfer knowledge from a high-resource modality to a low-resource one, a method that has since been widely adopted in various studies. Figure 5 illustrates the overall concept of cross-modal distillation and Table IV summarizes existing methods.
The early approach for CMKD was to use two parallel streams for two modalities during training, followed by a single stream during inference to produce feature maps similar to the teacher’s modality.[201, 202] applied this method to generate depth map from RGB. Later works focused on improving the imitation of the student’s output to match the teacher’s and reducing the gap between the modalities [203], with [204] employing two student networks, each attempting to replicate the other’s output. [205] adopted bidirectional distillation between events and frames.
The model proposed in [206] learned to generate 3D features from paired 2D inputs during the training phase and used them during inference. The structures in [207, 208, 209, 210] consisted of separate encoders and decoders to extract features from LiDAR and RGB inputs, aiming to align the different features from the two modalities and distill spatial knowledge from LiDAR to RGB. [211] used the same technique, replacing LiDAR with Radar, which provided lower spatial information, alongside RGB. Instead of using voxels for alignment, [212] generated depth and semantic features from both networks and aimed to make them similar. [213] employed a shared model to extract features from Thermal and RGB inputs in the student network, while using two heavier models for the teacher. The method proposed in [214] decomposed the process into task-specific components and enabled the distillation of knowledge from optical flow effects on the right part of the network.
To distill knowledge between text and image modalities, [215] extracted features separately and trained them on aligned instances, using soft alignment to supervise the student network. [216] applied masking techniques and used text features to guide the network in predicting masked images. [217] masked audio and video frames, attempting to reconstruct them to achieve a good representation for both modalities and align the two features effectively.
TABLE IV: Summary of Cross-modal Distillation Methods.
@@ IV-D Cross-modal Distillation
Gupta et al.[200] (2016) RGB Depth/Optical Flow
Garcia et al.[202] (2019) RGB + Depth RGB
EvDistill[205] (2021) RGB + Event Event
Hong et al.[207] (2022) DistillBEV[209] (2023) RGB + LiDAR RGB
UniDistill[208] (2023) RGB + LiDAR LiDAR
Lee et al.[214] (2023) RGB + Optical Flow RGB
Radocc[212] (2024) Multi-view + LiDAR Mult-view
CRKD[211] (2024) Multi-view + LiDAR Multi-view + Radar
Andonian et al. [215] (2022) CM-MaskSD [216] (2024) RGB + Text RGB + Text
XKD[217] (2024) Video RGB
IV-E Graph-based Distillation
@@
Graph-based knowledge distillation extends traditional KD methods by incorporating higher-order relational information between features or instances. Unlike approaches that primarily rely on pairwise similarity, graph-based methods utilize graph structures as a generic framework to capture intra-dependent relationships. In such methods, features or instances are represented as vertices, while the dependencies between them are modeled through edge weights. These graph structures encapsulate both local and global relationships, enabling the transfer of richer and more structured knowledge from teacher models to student models [85, 47, 45, 25, 86, 46, 218, 219, 220, 221, 222, 223]. Furthermore, graph structures can be effectively employed to model the flow of information during distillation, particularly in scenarios involving multiple teacher models, by regulating the knowledge transfer process [224, 225].
@@
In recent years, the advent of Graph Neural Networks (GNNs) has garnered significant attention, particularly in the fields of data mining and knowledge graph modeling. GNN distillation is primarily utilized to achieve two main objectives [226]: enhancing the performance of the original teacher model (performance improvement) [227, 228, 229, 230, 231, 232, 233, 234, 235] or creating a lightweight version of the GNN (compression). The latter involves distilling a larger and more complex GNN into a compact and efficient GNN [236, 237, 238, 239] or even an MLP [240, 241, 242, 243], making the distilled model suitable for real-time applications. To accomplish these goals, the information transferred between the teacher and the student typically includes logits [231, 232, 228, 227, 241, 240, 243, 236, 238, 239], feature representations [234, 232, 230, 227, 238], and the graph structure itself [235, 230, 229, 228, 242, 243, 237], which encapsulates the connectivity between the elements of the graph. These components collectively enable an effective transfer of structured knowledge, ensuring that the distilled model retains essential characteristics while meeting its respective performance or efficiency objectives.
IV-F Adaptive Distillation
@@
Adaptive distillation has recently garnered increased attention. By dynamically adjusting the parameters of the distillation process instead of maintaining a constant setup, the effectiveness of knowledge transfer can be enhanced. Table V summarizes adaptive distillation methods in different categories.
@@
Adaptive distillation can manifest in various levels and forms, such as adaptive loss definition [244, 49, 245], adaptive loss weighting [88, 246, 57], adaptive teacher pruning [18, 247, 248, 219, 249], and dynamically weighting different aspects of distillation based on sample importance [250, 251, 252, 253, 254, 255].
@@
One primary concern is that the teacher network may produce incorrect outputs that should not be directly transferred to the student. To address this issue, some approaches have introduced adaptive losses. [244] suggests an adaptive Cross Entropy loss that replaces incorrect probability maps with ground truth labels. [49] proposes a calibrated mask to prevent the teacher model’s erroneous representations from interfering with the student model’s training, while [245] integrates the positive prediction distribution of two teacher networks based on their correctness and cross-entropy magnitudes to provide a more accurate output distribution for guiding the student network.
@@
Moreover, certain studies have delved into the effects of altering the distillation loss throughout the distillation process. [88] was among the first to explore strategies where reducing the influence of the teacher network on the student can enhance performance. [246] indicates that the impact of the distillation loss should diminish as training progresses, with the student gradually assuming control of the training process towards the end. [57] introduces a weight decaying process to merge similarity distillation with vanilla distillation.
@@
Given the capacity disparity between teacher and student models, some works aim to narrow this gap. A notable example is TAKD [18], which employs a teaching assistant network as an intermediary teacher to facilitate better knowledge comprehension by the student model. Similarly, [219] utilizes an auxiliary teacher model before distillation, while [248] and [249] advocate for channel and feature pruning, respectively, before distillation. [247] implements filter pruning to reduce channels and applies a curriculum learning strategy to distill layers from easy to challenging levels.
@@
Lastly, certain methodologies adaptively conduct distillation based on the significance of a region or sample. [251] extracts detailed context-specific information from each training sample, [253] adaptively identifies confusing samples, and [254] adjusts the temperature parameter based on each sample’s energy level. [250] incorporates adaptive weights for distilling each mask in feature distillation, while [255] employs ground truth information to mask crucial regions for logit distillation. [252] calculates channel divergences between the teacher and student networks, converting them into distillation weights, and [256] assigns greater weight to losses in layers where training discrepancies between the teacher and student models are more pronounced during distillation.
TABLE V: Summary of adaptive distillation methods.
@@
Adaptive Loss Definition Park et al.[244], Wang et al.[49], CAG-DAKD [245]
Adaptive Loss Scaling Cho et al.[88], Zhou et al.[246], AICSD[57]
Adaptive Teacher Pruning TAKD[18], InDistill[247], Passalis et al.[219], PCD[248], AFMPM[249]
Adaptive Sample Importance MasKD[250], APD[251], HWD[252], CS-KD [253], Energy KD[254], Gou et al.[255]
IV-G Contrastive Distillation
Contrastive learning seeks to learn representations by contrasting similar (positive) and dissimilar (negative) samples in the feature space [257]. In KD, most research employs feature-based contrastive learning approaches, where anchor, positive, and negative sets are defined using features extracted by teacher and student models as shown in Figure 6. The contrastive distillation loss can be formulated as

### v2

@@ 4.2.3 Adversarial Robust Knowledge Distillation
As previously discussed, KD transfers knowledge from a large, pre-trained teacher network to a smaller student network, yet it does not inherently guarantee robustness against adversarial attacks. Adversarial Robust Distillation (ARD) (Goldblum et al., 2020) addresses this limitation by combining adversarial training with distillation. In ARD, the objective is to minimize the discrepancy between the student’s predictions and the teacher’s predictions under adversarial perturbations, defined as follows
@@
Building upon ARD, several approaches have been developed to enhance the transfer of robust knowledge. One group refines teacher prediction reliability: RSLAD (Zi et al., 2021) replaces the classification loss in eq.˜6 with a KL loss between teacher and student predictions to capitalize on robust soft labels from adversarially trained teachers, while introspective adversarial distillation (IAD)(Zhu et al., 2021a) and MTARD(Zhao et al., 2022e) further allow selective trust in teacher outputs and employ dual teachers for clean and adversarial scenarios, respectively. Another group enhances alignment and adaptability by synchronizing gradients and adaptively deriving perturbations through IGDM (Lee et al., 2023a) and AdaAD (Huang et al., 2023a); additional refinements are achieved in SmaraAD (Yin et al., 2024b) and IGKD-BML (Wang et al., 2021b) via Spearman Correlation, Class Activation Mapping, and attention-guided distillation to better emulate the teacher’s decision-making process.
@@
Advanced strategies tackle limitations in robustness transfer by introducing novel training schemes. DARWIN (Dong et al., 2024b) incorporates intermediate adversarial samples along with a triplet-based loss to balance natural, adversarial, and intermediate distributions, while PeerAiD (Jung et al., 2024) trains a peer model alongside the student for dynamic adaptation. Furthermore, DGAD (Park & Min, 2024) partitions the dataset into standard and adversarial distillation groups with consistency regularization to manage data imbalance, and both STARSHIP (Dong et al., 2024a) and TALD (Nguyen-Duc et al., 2023) further bolster robustness by transferring statistical attributes and sampling diverse adversarial examples via a Teacher Adversarial Local Distribution using Stein Variational Gradient Descent. These grouped techniques collectively advance the robustness and overall performance of student models in adversarial environments.
4.3 Multi-teacher Distillation
@@
In contrast to typical distillation scenarios where the student is trained using a single teacher, multi-teacher algorithms aim to train the student network by amalgamating knowledge from multiple teachers. This approach has the potential to enhance the student’s performance by leveraging an ensemble of teachers with diverse knowledge, thereby bolstering the student’s generalization capabilities. This concept was initially proposed in Hinton (2015), which utilized the average of logits from multiple teachers as softened labels. Subsequent works have explored variations on this theme: You et al. (2017) employed a voting strategy, while Sau & Balasubramanian (2016) recommended perturbing the teacher’s logits by adding noise, treating these perturbed outputs as distinct teachers that can act as a regularizer. TeKAP (Hossain et al., 2025) proposed an augmentation technique that generates multiple synthetic teacher knowledge by perturbing the knowledge of a single pretrained teacher. Fukuda et al. (2017) utilized a pool of teachers, randomly selecting a teacher for distillation each time. Papernot et al. (2016a) trained multiple teachers, with each focusing on a subset of the dataset, and employed a voting system to distill the knowledge of each expert teacher. Iordache et al. (2025) distilled knowledge from multiple teachers trained on distinct datasets. Furlanello et al. (2018) adopted a step-by-step strategy in which the student, at each stage, served as a teacher for the student in the subsequent step.
@@
In recent years, a variety of multi-teacher methods have been introduced, each addressing distinct issues. Managing the knowledge aggregation adaptively poses a challenge due to the capacity gap between the student and each teacher (Li & Bilen, 2020; Wang et al., 2023e; Zhang et al., 2023a; Shang et al., 2023b; Lin et al., 2024d; Cheng et al., 2024b; Wang & Xu, 2024; Li et al., 2024e). Lin et al. (2024d) employs adaptive temperature and a diverse aggregation strategy to enhance distillation performance, while Cheng et al. (2024b) separates the vanilla KD loss and assigns adaptive weights to each teacher based on the entropy of their predictions. Wang & Xu (2024) leverages data relation knowledge to dynamically allocate weights to teachers, and Zhang et al. (2023a) employs a meta-network for weighting each teacher effectively.
@@
Other methodologies have been proposed to effectively aggregate the knowledge of teachers for distillation (Luo et al., 2019; Chen et al., 2019b; Park & Kwak, 2020; Asif et al., 2020; Cao et al., 2023). Chen et al. (2019b) trains two distinct teachers, one for distilling intricate features and another for transferring general decision features. Park & Kwak (2020) and Asif et al. (2020) introduce additional branches to the student network for learning the features of teachers, while Cao et al. (2023) suggests a progressive approach for distilling knowledge from multiple teachers. Another significant research avenue in multi-teacher distillation involves using ensemble pre-trained teachers, referred to as Knowledge Amalgamation (KA) (Shen et al., 2019a; b; Vongkulbhisal et al., 2019; Luo et al., 2020; Amirkhani et al., 2021; Xu et al., 2022b; Zhang et al., 2023c; Li et al., 2024b). Through KA, publicly available trained networks can be repurposed in the context of multi-teacher distillation. Some approaches involve employing multiple teachers, each trained on different datasets (Dong et al., 2024b) or tasks (Shen et al., 2019a; Luo et al., 2019). Shen et al. (2019b) initially extracts task-specific knowledge from heterogeneous teachers sharing the same sub-task and then combines this extracted knowledge to construct the student network. Li et al. (2024b) proposes a KA framework based on uncertainty suppression, and Vongkulbhisal et al. (2019) suggests training a unified classifier from pre-trained classifiers from each source along with an unlabeled set of generic data.
@@
4.4 Cross-modal Distillation
In many scenarios, some modalities have rich annotations, while others lack sufficient labels. Cross-Modality Knowledge Distillation (CMKD) addresses this challenge by leveraging a well-annotated modality to improve the learning of a less-annotated one. This technique enables the model to benefit from the additional modality during training while remaining functional, even when that modality is unavailable at inference time. Gupta et al. (2016) introduced an innovative framework to transfer knowledge from a high-resource modality to a low-resource one, a method that has since been widely adopted in various studies. Figure 5 illustrates the overall concept of cross-modal distillation and Table 4 summarizes existing methods.
The early approach for CMKD was to use two parallel streams for two modalities during training, followed by a single stream during inference to produce feature maps similar to the teacher’s modality.Garcia et al. (2018; 2019) applied this method to generate depth map from RGB. Later works focused on improving the imitation of the student’s output to match the teacher’s and reducing the gap between the modalities (Huo et al., 2024), with Thoker & Gall (2019) employing two student networks, each attempting to replicate the other’s output. Wang et al. (2021d) adopted bidirectional distillation between events and frames.
The model proposed in Liu et al. (2021g) learned to generate 3D features from paired 2D inputs during the training phase and used them during inference. The structures in Hong et al. (2022b); Zhou et al. (2023b); Wang et al. (2023p); Klingner et al. (2023) consisted of separate encoders and decoders to extract features from LiDAR and RGB inputs, aiming to align the different features from the two modalities and distill spatial knowledge from LiDAR to RGB. Zhao et al. (2024b) used the same technique, replacing LiDAR with Radar, which provided lower spatial information, alongside RGB. Instead of using voxels for alignment, Zhang et al. (2024c) generated depth and semantic features from both networks and aimed to make them similar. Zhang et al. (2023i) employed a shared model to extract features from Thermal and RGB inputs in the student network, while using two heavier models for the teacher. The method proposed in Lee et al. (2023b) decomposed the process into task-specific components and enabled the distillation of knowledge from optical flow effects on the right part of the network.
To distill knowledge between text and image modalities, Andonian et al. (2022) extracted features separately and trained them on aligned instances, using soft alignment to supervise the student network. Wang et al. (2024j) applied masking techniques and used text features to guide the network in predicting masked images. Sarkar & Etemad (2024) masked audio and video frames, attempting to reconstruct them to achieve a good representation for both modalities and align the two features effectively.
Table 4: Summary of Cross-modal Distillation Methods.
@@ 4.4 Cross-modal Distillation
Gupta et al.(Gupta et al., 2016) RGB Depth/Optical Flow
Garcia et al.(Garcia et al., 2019) RGB + Depth RGB
EvDistill(Wang et al., 2021d) RGB + Event Event
Hong et al.(Hong et al., 2022b) DistillBEV(Wang et al., 2023p) RGB + LiDAR RGB
UniDistill(Zhou et al., 2023b) RGB + LiDAR LiDAR
Lee et al.(Lee et al., 2023b) RGB + Optical Flow RGB
Radocc(Zhang et al., 2024c) Multi-view + LiDAR Mult-view
CRKD(Zhao et al., 2024b) Multi-view + LiDAR Multi-view + Radar
Andonian et al. (Andonian et al., 2022) CM-MaskSD (Wang et al., 2024j) RGB + Text RGB + Text
XKD(Sarkar & Etemad, 2024) Video RGB
4.5 Graph-based Distillation
@@
Graph-based knowledge distillation extends traditional KD methods by incorporating higher-order relational information between features or instances. Unlike approaches that primarily rely on pairwise similarity, graph-based methods utilize graph structures as a generic framework to capture intra-dependent relationships. In such methods, features or instances are represented as vertices, while the dependencies between them are modeled through edge weights. These graph structures encapsulate both local and global relationships, enabling the transfer of richer and more structured knowledge from teacher models to student models (Zhang & Peng, 2018; Chen et al., 2020b; Park et al., 2019; Tung & Mori, 2019; Peng et al., 2019; Liu et al., 2019f; Lee & Song, 2019; Passalis et al., 2020; Zhou et al., 2021a; Chen et al., 2021e; Ghorbani et al., 2021; Lee & Song, 2021). Furthermore, graph structures can be effectively employed to model the flow of information during distillation, particularly in scenarios involving multiple teacher models, by regulating the knowledge transfer process (Luo et al., 2018; Minami et al., 2020).
@@
In recent years, the advent of Graph Neural Networks (GNNs) has garnered significant attention, particularly in the fields of data mining and knowledge graph modeling. GNN distillation is primarily utilized to achieve two main objectives (Tian et al., 2023a): enhancing the performance of the original teacher model (performance improvement) (Zhang et al., 2020a; 2021e; Chen et al., 2020e; Rezayi et al., 2021; Guo et al., 2023c; Huo et al., 2023; Zhu et al., 2024b; Yu et al., 2022a; Guo et al., 2022) or creating a lightweight version of the GNN (compression). The latter involves distilling a larger and more complex GNN into a compact and efficient GNN (Dong et al., 2023c; Yang et al., 2022b; He et al., 2022; Deng & Zhang, 2021) or even an MLP (Yang et al., 2021; Zhang et al., 2021d; Zheng et al., 2021; Tian et al., 2022a), making the distilled model suitable for real-time applications. To accomplish these goals, the information transferred between the teacher and the student typically includes logits (Guo et al., 2023c; Huo et al., 2023; Zhang et al., 2021e; 2020a; d; Yang et al., 2021; Tian et al., 2022a; Dong et al., 2023c; He et al., 2022; Deng & Zhang, 2021), feature representations (Yu et al., 2022a; Huo et al., 2023; Rezayi et al., 2021; Zhang et al., 2020a; He et al., 2022), and the graph structure itself (Guo et al., 2022; Rezayi et al., 2021; Chen et al., 2020e; Zhang et al., 2021e; Zheng et al., 2021; Tian et al., 2022a; Yang et al., 2022b), which encapsulates the connectivity between the elements of the graph. These components collectively enable an effective transfer of structured knowledge, ensuring that the distilled model retains essential characteristics while meeting its respective performance or efficiency objectives.
4.6 Adaptive Distillation
@@
Adaptive distillation has recently garnered increased attention. By dynamically adjusting the parameters of the distillation process instead of maintaining a constant setup, the effectiveness of knowledge transfer can be enhanced. Table 5 summarizes adaptive distillation methods in different categories.
@@
Adaptive distillation can manifest in various levels and forms, such as adaptive loss definition (Park & Heo, 2020; Wang et al., 2024g; Ma et al., 2024), adaptive loss weighting (Cho & Hariharan, 2019; Zhou et al., 2020; Mansourian et al., 2025), adaptive teacher pruning (Mirzadeh et al., 2020; Sarridis et al., 2025; Lee et al., 2024; Passalis et al., 2020; Guo et al., 2024a), and dynamically weighting different aspects of distillation based on sample importance (Huang et al., 2022b; Tian et al., 2022c; Sun et al., 2023b; 2024b; Kim et al., 2024c; Gou et al., 2024; Huang et al., 2025b).
@@
One primary concern is that the teacher network may produce incorrect outputs that should not be directly transferred to the student. To address this issue, some approaches have introduced adaptive losses. Park & Heo (2020) suggests an adaptive Cross Entropy loss that replaces incorrect probability maps with ground truth labels. Wang et al. (2024g) proposes a calibrated mask to prevent the teacher model’s erroneous representations from interfering with the student model’s training, while Ma et al. (2024) integrates the positive prediction distribution of two teacher networks based on their correctness and cross-entropy magnitudes to provide a more accurate output distribution for guiding the student network.
@@
Moreover, certain studies have delved into the effects of altering the distillation loss throughout the distillation process. Cho & Hariharan (2019) was among the first to explore strategies where reducing the influence of the teacher network on the student can enhance performance. Zhou et al. (2020) indicates that the impact of the distillation loss should diminish as training progresses, with the student gradually assuming control of the training process towards the end. Mansourian et al. (2025) introduces a weight decaying process to merge similarity distillation with vanilla distillation.
@@
Given the capacity disparity between teacher and student models, some works aim to narrow this gap. A notable example is TAKD (Mirzadeh et al., 2020), which employs a teaching assistant network as an intermediary teacher to facilitate better knowledge comprehension by the student model. Similarly, Passalis et al. (2020) utilizes an auxiliary teacher model before distillation, while Lee et al. (2024) and Guo et al. (2024a) advocate for channel and feature pruning, respectively, before distillation. Sarridis et al. (2025) implements filter pruning to reduce channels and applies a curriculum learning strategy to distill layers from easy to challenging levels.
@@
Lastly, certain methodologies adaptively conduct distillation based on the significance of a region or sample. Tian et al. (2022c) extracts detailed context-specific information from each training sample, Sun et al. (2024b) adaptively identifies confusing samples, and Kim et al. (2024c) adjusts the temperature parameter based on each sample’s energy level. Huang et al. (2022b) incorporates adaptive weights for distilling each mask in feature distillation, while Gou et al. (2024) employs ground truth information to mask crucial regions for logit distillation. Sun et al. (2023b) calculates channel divergences between the teacher and student networks, converting them into distillation weights, and Kang et al. (2024) assigns greater weight to losses in layers where training discrepancies between the teacher and student models are more pronounced during distillation. CALD (Huang et al., 2025b) leverages a meta-network to generate class-adaptive weights for distillation.
Table 5: Summary of adaptive distillation methods.
@@
Adaptive Loss Definition Park et al.(Park & Heo, 2020), Wang et al.(Wang et al., 2024g), CAG-DAKD (Ma et al., 2024), Dist+ (Huang et al., 2025a)
Adaptive Loss Scaling Cho et al.(Cho & Hariharan, 2019), Zhou et al.(Zhou et al., 2020), AICSD(Mansourian et al., 2025)
Adaptive Teacher Pruning TAKD(Mirzadeh et al., 2020), InDistill(Sarridis et al., 2025), Passalis et al.(Passalis et al., 2020), PCD(Lee et al., 2024), AFMPM(Guo et al., 2024a), ADG-KD(Li et al., 2025b)
Adaptive Sample Importance MasKD(Huang et al., 2022b), APD(Tian et al., 2022c), HWD(Sun et al., 2023b), CS-KD (Sun et al., 2024b), Energy KD(Kim et al., 2024c), Gou et al.(Gou et al., 2024), CALD(Huang et al., 2025b)
4.7 Contrastive Distillation
Contrastive learning seeks to learn representations by contrasting similar (positive) and dissimilar (negative) samples in the feature space (Hadsell et al., 2006). In KD, most research employs feature-based contrastive learning approaches, where anchor, positive, and negative sets are defined using features extracted by teacher and student models as shown in Figure 6. The contrastive distillation loss can be formulated as